# Why Fine-Tuning Struggles with Forgetting in Machine Unlearning? Theoretical Insights and a Remedial Approach

### v1

@@ 6 Experiment
In this section, we verify our theoretical insights by evaluating the effectiveness of the regularization-based FT machine unlearning methods through numerical experiments.
6.1 Experiment Setups
@@
Datasets and Models. The baseline method is the naive fine-tuning approach [8, 30] implemented on ResNet-18 [12] and we also include the golden retrained model for comparison. Our experiments will focus on image classification using the CIFAR-10 [15], CIFAR-100 [15], and SVHN [20] datasets. More details on the experimental setup will be provided in Appendix A.
@@
  • Remaining accuracy (RA): The accuracy of the model on the remaining dataset $D_{r}$ after unlearning, measuring how well the model retains its performance from pretrained model.
@@
6.2 Experiment Results
Table 2: Cifar-10, Cifar-100, SVHN Class-wise Forgetting. This table presents the performance of different unlearning methods, including FT, Sparse-FT, CE-FT, ICE-FT, and KL-FT, across CIFAR-10, CIFAR-100, and SVHN datasets. The evaluation metrics include Unlearning Accuracy (UA), MIA-Efficacy, Retaining Accuracy (RA), Test Accuracy (TA), and RunTime. Results are reported as mean $\pm$ standard deviation over five independent trials. A performance gap is computed against the retrain method, showing how each approach performs relative to the golden retraining method.
Cifar-10 Class-wise Forgetting
Methods UA MIA-Efficacy RA TA Run Time
Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.87_{\pm 0.14}$ $77.00$
FT $20.89_{\pm 4.12}(79.11)$ $74.32_{\pm 11.90}(25.68)$ $99.76_{\pm 0.12}(0.24)$ $93.73_{\pm 0.22}(1.14)$ $4.48$
CE-FT $100.00_{\pm 0.00}(0.00)$ $100.00_{\pm 0.00}(0.00)$ $75.76_{\pm 5.03}(24.24)$ $68.67_{\pm 5.11}(26.20)$ $6.49$
ICE-FT $100.00_{\pm 0.00}(0.00)$ $100.00_{\pm 0.00}(0.00)$ $92.22_{\pm 0.72}(7.78)$ $84.22_{\pm 1.12}(10.65)$ $6.43$
KL-FT $99.17_{\pm 0.29}(0.83)$ $100.00_{\pm 0.00}(0.00)$ $99.06_{\pm 0.51}(0.94)$ $92.54_{\pm 0.67}(2.33)$ $5.50$
Cifar-100 Class-wise Forgetting
Methods UA MIA-Efficacy RA TA Run Time
Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $99.81_{\pm 0.06}$ $75.14_{\pm 0.12}$ $81.00$
FT $34.44_{\pm 19.89}(65.56)$ $87.64_{\pm 10.31}(12.36)$ $99.79_{\pm 0.10}(0.21)$ $75.07_{\pm 0.56}(0.07)$ $5.20$
CE-FT $99.87_{\pm 0.13}(0.13)$ $99.95_{\pm 0.05}(0.05)$ $94.71_{\pm 1.01}(5.10)$ $63.64_{\pm 0.93}(11.50)$ $6.23$
ICE-FT $100.00_{\pm 0.00}(0.00)$ $100.00_{\pm 0.00}(0.00)$ $96.66_{\pm 1.28}(3.15)$ $66.65_{\pm 2.08}(8.49)$ $4.31$
KL-FT $95.20_{\pm 2.31}(4.80)$ $100.00_{\pm 0.00}(0.00)$ $99.26_{\pm 0.16}(0.55)$ $73.11_{\pm 0.42}(2.03)$ $6.20$
SVHN Class-wise Forgetting
Methods UA MIA-Efficacy RA TA Run Time
Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $88.20_{\pm 0.75}$ $72.00$
FT $17.47_{\pm 7.29}(82.53)$ $99.95_{\pm 0.05}(0.05)$ $100_{\pm 0.00}(0.00)$ $93.55_{\pm 0.63}(5.35)$ $5.01$
CE-FT $100_{\pm 0.00}(0.00)$ $100_{\pm 0.00}(0.00)$ $97.27_{\pm 2.70}(3.73)$ $79.77_{\pm 3.51}(10.87)$ $5.48$
ICE-FT $100.00_{\pm 0.00}(0.00)$ $100.00_{\pm 0.00}(0.00)$ $99.99_{\pm 0.01}(0.01)$ $85.56_{\pm 0.32}(2.65)$ $4.48$
KL-FT $97.24_{\pm 0.90}(2.76)$ $100.00_{\pm 0.00}(0.00)$ $99.95_{\pm 0.05}(0.05)$ $87.54_{\pm 0.15}(0.66)$ $5.23$
Performance Comparison among Regularization-Based Fine-Tuning Methods. In Table 2, we explore the impact of different regularization terms on the performance of various FT-based methods. It is evident that the regularization term consistently enhances the model’s unlearning accuracy. Specifically, in the CIFAR-10 experiments, all regularization-based methods show a significant improvement in UA compared to the baseline FT ($20.89\%$). Both ICE-FT and CE-FT achieve perfect performance in UA and MIA-Efficacy, showing no gap with the golden model. However, it is also noticeable that their RA and TA, especially for the CE-FT method, decrease dramatically, indicating that the improvement in UA comes at the expense of reduced RA and TA. Notably, our KL-FT method achieves the most comparable RA of $99.06\%$ and TA of $92.54\%$ relative to the baseline Retrain ($100.00\%$ and $94.87\%$) and FT ($99.76\%$ and $93.73\%$), without the extreme RA decline seen in other methods. Similarly, in the CIFAR-100 experiments, KL-FT achieves the best RA of $99.26\%$ and TA of $73.11\%$, outperforming all other methods, with only a minor decline in UA. These highlight the effectiveness of KL-FT in balancing unlearning and remaining accuracy. For the SVHN dataset, the FT method shows even more challenges in forgetting classes, with a UA of $17.47\%$. In contrast, all regularization-based methods, including CE-FT, ICE-FT, KL-FT, achieve perfect UA, signifying complete forgetting. ICE-FT achieves the best TA of $85.56\%$, closely followed by KL-FT with a TA of $87.54\%$.
(a) KL
(b) CE vs ICE
(c) RA of three methods
Figure 4: This figure shows the impact of varying the regularization parameter $\alpha$ on the accuracy metrics for KL-FT, CE-FT, and ICE-FT. Figure 4(a) illustrates the sensitivity of the KL method across various evaluation metrics within the range of 0 to 1. Figure 4(b) shows the impact of different regularization focuses by comparing the performance of CE-FT and ICE-FT. Figure 4(c) presents the RA behavior for all three methods (KL, CE, and ICE-FT).
ICE-FT builds upon the CE-FT by adjusting the focus of regularization. Instead of prioritizing forgetting during the unlearning process, ICE-FT balances the two, placing more emphasis on RA. This change allows ICE-FT to achieve perfect unlearning (UA of 100.00%) while maintaining significantly higher RA and TA compared to CE-FT across all datasets. These results align with our previous theoretical analysis, which suggests that focusing on retention does not significantly impact UA, but shifting the focus away from retention compromises RA.
Sensitivity of regularization parameter $\alpha$. The regularization parameter $\alpha$ is a crucial hyperparameter in our proposed method. To demonstrate its effect, we conduct numerical experiments on the CIFAR-10 dataset, showing how the regularization parameter impacts the unlearning performance of various regularization-based FT methods. We first examine the sensitivity of the KL method across various evaluation metrics within the range of $(0,1]$. As shown in Figure 4(a), as $\alpha$ increases, the RA gradually declines from near-perfect performance ( $\sim 100\%$ ) down to $\sim 94\%$ at $\alpha=0.8$, indicating that stronger regularization negatively affects the retention of information. Meanwhile, the Test Accuracy follows the same downward trend as the Remaining Accuracy, dropping from $\sim 94\%$ to $\sim 88\%$. Additionally, we explore how different regularization focuses impact unlearning performance. In Figure 4(b), we observe that the RA of CE-FT decreases gradually with increasing $\alpha$, and UA rises slightly before stabilizing, the MIA-Efficacy stays consistently high. In contrast, the ICE-FT method prioritizes retaining accuracy on the remaining data, resulting in a higher Retain Accuracy than CE-FT across all $\alpha$ values. The UA slightly decreases but remains competitive, while MIA-Efficacy and Test Accuracy follow similar trends to those seen in CE-FT. This suggests that ICE-FT achieves a better balance between unlearning and accuracy retention than standard CE-FT, which aligns with our previous analysis. Finally, we illustrate the RA and TA behavior for all three methods in Figure 4(c).
Discriminative Regularization beyond FT-based methods. While our study focuses on the use of discriminative regularization within fine-tuning (FT)-based machine unlearning methods, the potential applications of this approach extend beyond FT. Our discriminative regularization framework, designed to balance unlearning accuracy (UA) and remaining accuracy (RA), can be adapted to other unlearning techniques, including Gradient Ascent, Fisher Forgetting, or Influence Unlearning. The flexibility of our discriminative regularization framework offers potential for improving a wide range of unlearning strategies, which we leave for future exploration.
@@ 7 Conclusion
In conclusion, we present the first theoretical analysis of fine-tuning methods for machine unlearning within a linear regression framework. Our analysis, covering two scenarios—distinct and overlapping feature sets—demonstrates that while fine-tuning can achieve optimal remaining accuracy (RA), it fails to fully unlearn the forgetting dataset. Our analysis on the failure of naive fine-tuning methods stems from the pretrained model’s retention of forgetting data, and we propose a theoretical approach to mitigate this issue. By introducing a discriminative regularization term, we effectively balance unlearning accuracy and remaining accuracy. Experimental validation on both synthetic and real-world datasets supports our theoretical insights, showing that regularization-based methods significantly improve unlearning performance without compromising retention.
@@ A.1 Verification via Simulation
Data Generation. We constructed two data matrices, $\mathbf{X}_{r}$ and $\mathbf{X}_{f}$, representing the remaining data and the forgetting data, respectively. The remaining data matrix, $\mathbf{X}_{r}^{\top}$, was structured as $[\mathbf{R}^{\top},\mathbf{L}_{1}^{\top},\mathbf{0}]$, and the forgetting data matrix, $\mathbf{X}_{f}^{\top}$, as $[\mathbf{0},\mathbf{L}_{2}^{\top},\mathbf{F}^{\top}]$, where $\mathbf{L}_{1}^{\top}=\mathbf{L}_{2}^{\top}=\mathbf{0}$ to enforce the non-overlapping case. Here, $\mathbf{R}^{\top}$ and $\mathbf{F}^{\top}$ are random matrices corresponding to different feature sets, and the zeros represent the distinct features across the datasets. We set the total number of data points to $n=40$ and the total number of features to $d=40$. The remaining data consisted of $n_{r}=30$ samples with $d_{r}=20$ features, while the forgetting data comprised $n_{f}=10$ samples with $d_{f}=d-d_{r}=20$ features. To simulate a controlled environment, we fixed the number of overlapping features to $d_{o}=0$ and $d_{o}=8$ for non-overlapping case and overlapping case, respectively.
@@
Model Training. To compare the effects of fine-tuning, we considered two models: the fine-tuning model $\mathbf{w}_{t}$ and the golden model $\mathbf{w}_{g}$. Specifically, $\mathbf{w}_{t}$ was obtained by fine-tuning on a subset of the remaining data, denoted as $\mathbf{X}_{t}$, which consisted of the first $n_{t}$ data points from $\mathbf{X}_{r}$. The value of $n_{t}$ varied from 1 to $n_{r}-1$ to study the impact of the fine-tuning data size. The initial model $\mathbf{w}_{o}$ was derived from the entire dataset $\mathbf{X}$ and calculated by the Equation 1. $\mathbf{w}_{g}$ was trained from scratch on the entire remaining data $\mathbf{X}_{r}$ and computed by solving Equation 2. If considering the regularization case in synthetic data, the regularized pretrained model will be constructed by zeroing out the coefficients corresponding to the forgetting data features with (without) overlapping features.
@@
  • Remaining Data Loss (RA Loss): $\operatorname{MSE}_{\mathrm{RA}}(\mathbf{w})=\frac{1}{n_{r}}\|\mathbf{X}_{r}\mathbf{w}-\mathbf{y}_{r}\|^{2}$.
@@
Experimental Results Figure 1(c) and Figure 1(d) illustrate that the regularized fine-tuning method discussed in Section 4 can significantly improve unlearning accuracy while preserving the remaining accuracy. Specifically, both the remaining loss and unlearning loss of $\hat{\mathbf{w}}_{t}$ perfectly match those of the golden model under both distinct and overlapping feature scenarios. Additionally, Figure 2 present comparisons of machine unlearning loss for different approaches to handling overlapping features: Figure 2(a) retains overlapping features from the pretrained model, demonstrating matching performance between the regularized $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; whereas Figure 2(b) discards the overlapping features, resulting in a decline in remaining accuracy. These empirical results align well with our theoretical findings.
A.2 Additional Real-world Experiments
Unlearning Setup The unlearning setup centers on the FT-based procedure. During training, the model is updated using the remaining dataset, while Kullback–Leibler divergence/Cross-entropy loss regularization is applied to the forget dataset to enforce unlearning. The corresponding regularization modifies the model’s predictions by encouraging it to generate incorrect outputs for the forgetting data. Specifically, KL divergence is computed between the model’s output and shifted incorrect labels, ensuring the model no longer retains knowledge of the forgetting data. Additionally, cross-entropy loss between the model’s output and the incorrect labels further supports the unlearning process. Throughout training, the optimizer updates the model based on the combined loss. Our experiments focus on class-wise forgetting, and we run the process 5 times, reporting the mean and standard deviation of the performances.
We summarize the datasets and model configurations in Tab. 3.
Table 3: Dataset and model setups.
Dataset CIFAR-10 SVHN CIFAR-100
Settings ResNet-18 ResNet-18 ResNet-18
Batch Size 256 256 256
Table 4: Comparison of CE and ICE results across various $\alpha$ values on Cifar-10 dataset.
$\alpha$ CE-FT ICE-FT
UA MIA RA Test UA MIA RA Test
0.1 $100.00$ $100.00$ $68.48$ $61.11$ $100.00$ $100.00$ $91.66$ $83.67$
0.2 $100.00$ $100.00$ $70.7$ $65.41$ $100.00$ $100.00$ $87.97$ $80.27$
0.3 $100.00$ $100.00$ $75.67$ $67.81$ $100.00$ $100.00$ $85.24$ $77.92$
0.4 $100.00$ $100.00$ $77.73$ $69.91$ $100.00$ $100.00$ $87.79$ $80.02$
0.5 $100.00$ $100.00$ $79.6$ $71.73$ $100.00$ $100.00$ $84.31$ $76.89$
0.6 $100.00$ $100.00$ $80.97$ $73.06$ $100.00$ $100.00$ $80.18$ $72.56$
0.7 $100.00$ $100.00$ $81.49$ $73.71$ $100.00$ $100.00$ $82.03$ $74.33$
0.8 $100.00$ $100.00$ $81.23$ $73.94$ $100.00$ $100.00$ $81.97$ $74.98$
0.9 $100.00$ $100.00$ $78.89$ $71.2$ $100.00$ $100.00$ $82.51$ $74.92$
@@ A.2 Additional Real-world Experiments
Figure 5: Visualization of Remaining Data and Forgetting Data Features Across Various Dataset. Figures 5(a)-5(c) focus on classes 3, 6, and 9 in CIFAR-10 and Figures 5(d)-5(f) focus on classes 30, 60, and 90 in CIFAR-100.
@@ B.1 Useful properties
Let $\mathbf{P}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Under 1, the remaining(forgetting) dataset matrix can be denoted as $\mathbf{X}_{r}=\Big{[}\begin{array}[]{l}\mathbf{R}\\ \mathbf{0}\end{array}\Big{]}$ and $\mathbf{X}_{f}=\Big{[}\begin{array}[]{l}\mathbf{0}\\ \mathbf{F}\end{array}\Big{]}$, where $\mathbf{R}\subseteq\mathbb{R}^{d_{r}\times(1-n_{f})}$ and $\mathbf{F}\subseteq\mathbb{R}^{d_{f}\times n_{f}}$ correspond to the non-zero parts. Then, it holds that:
@@
Firstly, based on the data composition, the overall dataset holds $\mathbf{X}=\Big{[}\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}\Big{]}$. Therefore, it follows:
$\mathbf{P}$ $=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}=\Big{[}\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}\Big{]}(\Big{[}\begin{array}[]{cc}\mathbf{R}^{\top}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}^{\top}\end{array}\Big{]}\Big{[}\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}\Big{]})^{-1}\Big{[}\begin{array}[]{cc}\mathbf{R}^{\top}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}^{\top}\end{array}\Big{]}$
$=\Big{[}\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}\Big{]}\Big{[}\begin{array}[]{cc}(\mathbf{R}^{\top}\mathbf{R})^{-1}&\mathbf{0}\\ \mathbf{0}&\Big{(}\mathbf{F}^{\top}\mathbf{F}\Big{)}^{-1}\end{array}\Big{]}\Big{[}\begin{array}[]{cc}\mathbf{R}^{\top}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}^{\top}\end{array}\Big{]}.$
@@
The proof is complete. ∎
Corollary 2 (Minimum Norm Solution).

### v2

@@ 5.2 Experiment
In the following, we verify our theoretical insights by evaluating the effectiveness of the mask-based FT machine unlearning methods through numerical experiments.
@@
Datasets and Models. Our baseline methods include the naive fine-tuning approach [8, 29] and SalUn [6], both implemented on ResNet-18 [12]. For comparison, we also include the golden retrained model as a reference. Our experiments will focus on image classification using the CIFAR-10 [15] and CIFAR-100 [15]. More details on the experimental setup will be provided in Appendix A.
@@
  • Retaining accuracy (RA): The accuracy of the model on the remaining dataset $D_{r}$ after unlearning, measuring how well the model retains its performance from the pretrained model.
@@
  • Average Disparity (Avg. Disparity): The mean absolute difference between the performance metrics of the unlearned model and the retrained model, calculated across all evaluation aspects.
@@
Masking Strategies Enhance Unlearning Accuracy in Naive Fine-Tuning. The performance of Retention-Based Masking (Ours) and Forgetting-Based Masking (SalUn) against Fine-Tuning highlights how masking significantly improves unlearning accuracy, measured as the distance from the retrained model. Specifically, in CIFAR-10 Class-wise Forgetting, naive fine-tuning exhibits a significant gap of $91.64$ compared to the retrained model, which is drastically reduced by SalUn ($0.22$) and further refined by RBM to $0.02$. Similarly, for CIFAR-100 Class-wise Forgetting, fine-tuning shows a large gap of $87.96$, reduced to $0.00$ by SalUn and maintained by RBM at $0.00$. In random forgetting scenarios, in CIFAR-10 Random Forgetting, fine-tuning achieves a gap of $5.62$, while SalUn reduces this to $3.42$ and RBM further minimizes it to $2.56$. A similar trend is observed for CIFAR-100 Random Forgetting, where fine-tuning starts with a gap of $24.64$, SalUn reduces it to $3.00$, and RBM achieves a further improvement to $1.22$. These results demonstrate that masking strategies, whether leveraging forgetting-based or retention-based saliency, are essential for enhancing unlearning accuracy in naive fine-tuning.
Retention-Based Masking Improves Retaining Accuracy. Next, we compare retention-based masking (Ours) and forgetting-based masking (SalUn) to evaluate how different masking approaches impact retaining accuracy, measured by the distance from the retrained model. For CIFAR-10 Random Forgetting, SalUn achieves a gap of $0.38$, while RBM improves it to $0.28$. For CIFAR-100 Random Forgetting, SalUn shows a gap of $1.33$, whereas RBM significantly reduces it to $1.24$. In class-wise forgetting scenarios, RBM consistently maintains an advantage: for CIFAR-10 Classwise Forgetting, SalUn achieves a gap of $0.53$ compared to $0.31$ for RBM. Similarly, for CIFAR-100 Class-wise Forgetting, SalUn achieves a gap of $0.35$, while RBM further improves this to $0.08$. These results underscore that the retention-based saliency approach is more effective in preserving the performance of the remaining dataset, aligning with our theoretical insights in Theorem 3.
Retention-Based Masking Minimizes Average Disparity Across Tasks. In addition to improving retaining accuracy, RBM also reduces average disparity across all tasks. For CIFAR-10 Random Forgetting, RBM reduces the disparity to $0.88$, compared to $1.39$ for SalUn. Similarly, in CIFAR-100 Random Forgetting, RBM achieves a reduced disparity of 6.80, lower than $7.91$ for SalUn. In CIFAR-10 Classwise Forgetting, RBM reduces the disparity to $0.42$, compared to $0.50$ for SalUn. For CIFAR-100 Class-wise Forgetting, RBM achieves an exceptionally low disparity of $0.19$, significantly lower than $0.35$ for SalUn. These results further validate the robustness of RBM in effectively balancing retaining and unlearning objectives.
@@ 6 Conclusion
In conclusion, we present the first theoretical analysis of fine-tuning methods for machine unlearning within a linear regression framework. Our analysis, covering two scenarios—distinct and overlapping feature sets—demonstrates that while fine-tuning can achieve optimal retaining accuracy, it fails to fully unlearn the forgetting dataset. This failure arises from the pretrained model retaining information about the forgetting data. To address this, we propose a theoretical solution and introduce Retention-Based Masking (RBM), a strategy that constructs masks based on the remaining dataset to preserve overlapping features. Our experimental results demonstrate that RBM achieves a better balance between unlearning and retaining objectives compared to existing masking approaches.
@@ A.1 Verification via Simulation
Data Generation. We constructed two data matrices, $\mathbf{X}_{r}$ and $\mathbf{X}_{f}$, representing the remaining data and the forgetting data, respectively. The remaining data matrix, $\mathbf{X}_{r}^{\top}$, was structured as $[\mathbf{R}^{\top},\mathbf{L}_{1}^{\top},\mathbf{0}]$, and the forgetting data matrix, $\mathbf{X}_{f}^{\top}$, as $[\mathbf{0},\mathbf{L}_{2}^{\top},\mathbf{F}^{\top}]$, where $\mathbf{L}_{1}^{\top}=\mathbf{L}_{2}^{\top}=\mathbf{0}$ to enforce the non-overlapping case. Here, $\mathbf{R}^{\top}$ and $\mathbf{F}^{\top}$ are random matrices corresponding to different feature sets, and the zeros represent the distinct features across the datasets. We set the total number of data points to $n=40$ and the total number of features to $d=40$. The remaining data consisted of $n_{r}=30$ samples with $d_{r}=20$ features, while the forgetting data comprised $n_{f}=10$ samples with $d_{f}=d-d_{r}=20$ features. To simulate a controlled environment, we fixed the number of overlapping features to $d_{lap}=0$ and $d_{lap}=8$ for non-overlapping case and overlapping case, respectively.
@@
Model Training. To compare the effects of fine-tuning, we considered two models: the fine-tuning model $\mathbf{w}_{t}$ and the golden model $\mathbf{w}_{g}$. Specifically, $\mathbf{w}_{t}$ was obtained by fine-tuning on a subset of the remaining data, denoted as $\mathbf{X}_{t}$, which consisted of the first $n_{t}$ data points from $\mathbf{X}_{r}$. The value of $n_{t}$ varied from 1 to $n_{r}-1$ to study the impact of the fine-tuning data size. The initial model $\mathbf{w}_{o}$ was derived from the entire dataset $\mathbf{X}$ and calculated by the Equation 1. $\mathbf{w}_{g}$ was trained from scratch on the entire remaining data $\mathbf{X}_{r}$ and computed by solving Equation 2. If considering the masked case in synthetic data, the masked pretrained model will be constructed by zeroing out the coefficients corresponding to the forgetting data features with (without) overlapping features.
@@
  • Remaining Data Loss (RA Loss):$\operatorname{MSE}_{\mathrm{RA}}(\mathbf{w})=\frac{1}{n_{r}}\|\mathbf{X}_{r}\mathbf{w}-\mathbf{y}_{r}\|^{2}$
@@
Experimental Results Figure 1(c) and Figure 1(d) illustrate that the masked fine-tuning method discussed in Section 4 can significantly improve unlearning accuracy while preserving the retaining accuracy. Specifically, both the remaining loss and unlearning loss of $\hat{\mathbf{w}}_{t}$ perfectly match those of the golden model under both distinct and overlapping feature scenarios. Additionally, Figure 2 present comparisons of machine unlearning loss for different approaches to handling overlapping features: Figure 2(a) retains overlapping features from the pretrained model, demonstrating matching performance between the masked $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; whereas Figure 2(b) discards the overlapping features, resulting in a decline in retaining accuracy. These empirical results align well with our theoretical findings.
A.2 Additional Real-world Details
Experimental Setup. We conduct our evaluations using the ResNet-18 backbone across the methods. The network is initially trained for classification over the CIFAR datasets for $182$ epochs with an initial learning rate of $0.1$ following a cosine decay schedule. For the unlearning procedure, the learning rate is set to $0.02$ for our method and $0.013$ for the SalUn method, following the recommendations from the official repository. We set the number of unlearning epochs to $10$. However, we observe that both methods benefit from a reduced number of unlearning epochs $(5)$ in the random forgetting scenario on CIFAR-100. We set the sparsity at $50\%$. Unlearning using the fine-tuning method employs a learning rate of $0.01$ for $10$ epochs. All methods utilize the SGD optimizer.
@@ A.2 Additional Real-world Details
Figure 3: Visualization of Remaining Data and Forgetting Data Features Across Various Dataset. Figures 3(a)-3(c) focus on classes 3, 6, and 9 in CIFAR-10 and Figures 3(d)-3(f) focus on classes 30, 60, and 90 in CIFAR-100.
@@ B.1 Useful properties
Let $\mathbf{P}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Under 1, the remaining(forgetting) dataset matrix can be denoted as $\mathbf{X}_{r}=[\begin{array}[]{l}\mathbf{R}\\ \mathbf{0}\end{array}]$ and $\mathbf{X}_{f}=[\begin{array}[]{l}\mathbf{0}\\ \mathbf{F}\end{array}]$, where $\mathbf{R}\subseteq\mathbb{R}^{d_{r}\times{(n-n_{f})}}$ and $\mathbf{F}\subseteq\mathbb{R}^{d_{f}\times n_{f}}$ correspond to the non-zero parts. Then, it holds that:
@@
Firstly, based on the data composition, the overall dataset holds $\mathbf{X}=[\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}]$. Therefore, it follows:
$\mathbf{P}$ $=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}=[\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}]([\begin{array}[]{cc}\mathbf{R}^{\top}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}^{\top}\end{array}][\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}])^{-1}[\begin{array}[]{cc}\mathbf{R}^{\top}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}^{\top}\end{array}]$
$=[\begin{array}[]{cc}\mathbf{R}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}\end{array}][\begin{array}[]{cc}(\mathbf{R}^{\top}\mathbf{R})^{-1}&\mathbf{0}\\ \mathbf{0}&(\mathbf{F}^{\top}\mathbf{F})^{-1}\end{array}][\begin{array}[]{cc}\mathbf{R}^{\top}&\mathbf{0}\\ \mathbf{0}&\mathbf{F}^{\top}\end{array}].$
@@
Corollary 2 (Minimum Norm Solution 1).